1. To find the radius of a cylinder, what information do you need to know? Why?
2. A cylindrical tank could hold 100.48m (raise to 3 ung m) of water. If the height of the tank is 2m, what is the radius of its base?
3. How many cubic meters of water a cylindrical tank can hold if its radius is 2m and its height is 3m?
1
Expert's answer
2012-02-02T11:13:16-0500
1. There is a formula that describe cylinder's volume, height and radius:
V = H*2*pi*R,

so R = V / (2pi*H), so we need volume and height of cylinder

2. V = 100.48 m^3

R = V / (2pi*H) = 100.48 / (2 * 3.14 * 2) = 8 m

3. V = H*2*pi*R = 3 * 2 * 3.14 * 2 = 37.68 m^3